# Situationally-Aware Dynamics Learning

### v2

@@ IV Method
We propose an online, unsupervised representation learning method to capture and adapt to unknown, variable ego- and environmental-factors. Our approach models online changes in the distribution governing the robot’s transition dynamics as it interacts with the world. These changes encapsulate the influence of latent, unobserved factors—such as unmeasured variables or emergent phenomena—that are challenging to predict or exhaustively anticipate. By leveraging the learned distribution of local dynamics, the robot refines its transition model, enabling more effective planning and control in downstream tasks.
IV-A Online Transition Distribution Modeling
@@
$=\frac{\Pr(S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime})}{\Pr(S_{t}=s,A_{t}=a)}.$ (8)
In particular, $P(s^{\prime}\mid s,a)\propto\Pr(S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime})$. However, as previously noted, in real-world applications $P(s^{\prime}\mid s,a)$ usually fails to accurately capture the true dynamics of the world. Generally, this shortfall arises from latent, unobserved factors—such as unmeasured variables or phenomena for which the robot lacks the necessary sensors—that significantly impact state transitions. Thus, our key insight is that by modeling online the joint transition distribution $\Pr(S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime})$ that best explains the UDGP governing the current state transitions, we can account for the latent factors that influence the world dynamics and consequently improve performance on downstream tasks.
To achieve this, we break the problem of online transition distribution modeling into determining local joint transition distributions. By “local”, we mean around the latest set or trajectory of states experienced by the agent. For example, an UGV that is stuck in a slippery area while climbing a mountain will estimate a different local joint transition distribution than the same robot when it is moving over a flat and smooth sidewalk. Figure 1 illustrates this concept, showing how the same action executed from an identical state can lead to vastly different outcomes whenever some important latent factors are not accounted for. Crucially, such discrepancies are often better attributed to an incomplete state representation—e.g., the robot lacks sensors, such as a camera, to detect terrain types—rather than inherent aleatoric errors [^1] While our approach does account for aleatoric uncertainty in the transition model, as done in traditional dynamics learning methods, our primary focus is on reducing epistemic uncertainty by identifying and accounting for the latent factors that currently influence the robot’s dynamics..
@@
However, to effectively model the local dynamics using Eq. (9), we need to estimate $\boldsymbol{\eta}=\{\boldsymbol{\mu},\Lambda\}$ online and simultaneously detect when a different $\boldsymbol{\eta}^{\prime}$ describes the current situation of the robot better. For this, we extend BOCD to a multivariate setting and apply it to discover online and unsupervisedly the multiple sets $\boldsymbol{\eta}_{\rho}$. We proceed to introduce our mutivariate extension of BOCD in Section IV-A1. Then, in Section IV-A2 we show how we can leverage this results to model online the joint transition distribution of the robot’s dynamics.
IV-A1 Multivariate BOCD:
@@
Since we want to understand when the parameters $\boldsymbol{\eta}$ of the UDGP have changed, we need to define a prior for them. For this, we propose the Normal-Wishart distribution. This conjugate prior is particularly useful because it allows for closed form belief updates for $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and $\Lambda\in\mathbb{R}^{d\times d}$ as more data becomes available. It is defined as follows [murphy2007conjugateforms]:
@@
Essentially, the Normal-Wishart prior combines two distributions: it leverages the Normal distribution for the mean $\boldsymbol{\mu}$, incorporating prior knowledge $\boldsymbol{\mu}_{0}$ about the likely location of $\boldsymbol{\mu}$, and it employs the Wishart distribution to model the precision matrix $\Lambda$, acting as the prior on the inverse covariance. The Wishart distribution is flexible, encoding varying levels of confidence in the scale of the precision matrix through its degrees of freedom parameter and scale matrix, $T$. Parameters $\kappa_{0}$ and $\nu_{0}$ allow us to control the strength of our prior beliefs, with $\kappa_{0}$ scaling the covariance of the mean and $\nu_{0}$ defining the degrees of freedom of the Wishart distribution. This prior choice is satisfactory for our application since it assumes an unknown mean vector $\boldsymbol{\mu}$ and precision matrix $\Lambda$, which enables flexibility by allowing uncertainty in the covariance structure to be encoded in a principled way [degroot1970optimal, murphy2007conjugateforms]. Consequently, it spares us from imposing restrictive assumptions on the covariance matrix, such as isotropy, homogeneity, or mean-dependence, making it highly suitable for modeling multivariate normal distributions with little assumptions about its parameters.
With the sampling model Eq. (10) and prior Eq. (11) determined, we proceed to make the connection to BOCD.
Specifically, we will determine the closed form solutions for the growth probability Eq. (5) and changepoint probability Eq. (6), from which we can recover the run length probability Eq. (7). We begin by assuming that the prior on the changepoint probability is $P_{\mathtt{gap}}(g)\sim\mathtt{geometric}(\lambda)$. This makes the process memoryless and the hazard function constant at $H(\tau)=1/\lambda$, where $\lambda$ can be interpreted as the expected time between changepoints [adams2007bayesian].
@@
The proof follows from the results of the posterior distribution parameters for the Normal-Wishart prior in [degroot1970optimal, Sec. 9.10]. Specifically, they show that for a set of samples $\{\boldsymbol{x}_{i}:\boldsymbol{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$, the posterior parameters for Eq. (11), given prior parameters $\{\boldsymbol{\mu}_{0},\kappa_{0},T_{0},\nu_{0}\}$, are:
@@
Then, the derivation follows from the fact that $n=1$ for the observation $\boldsymbol{x}_{t}$ at time $t$, and the fact that the parameters $\boldsymbol{\eta}_{t-1}^{(r)}=\{\boldsymbol{\mu}_{t-1}^{(r)},T_{t-1}^{(r)},\nu_{t-1}^{(r)},\kappa_{t-1}^{(r)}\}$ from the previous time step are the current prior. ∎
@@
From the BOCD framework, we assumed it is possible to compute the predictive distribution conditional on a given run length [adams2007bayesian]:
@@
Additionally, $P(\boldsymbol{x}_{t+1}|\boldsymbol{x}_{1:t})$ depends only on the data of the current run $\boldsymbol{x}_{t}^{(r)}$. Thus, $P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{1:t})=P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{t}^{(r)})$. Therefore, using the result from [murphy2007conjugateforms] that says that the posterior predictive of the Normal-Wishart prior is:
@@
IV-A2 Online Situation Modeling:
To determine when $T_{L}$ changes and given that we are modeling the system’s situations via distribution parameters $\boldsymbol{\eta}_{0}\coloneqq\{\boldsymbol{\mu}_{0},\kappa_{0},T_{0},\nu_{0}\}$, we can use our previous multivariate extension to BOCD. Specifically, we do so as detailed in Algorithm 2. Concretely, we first determine if we have changed situations (Lemma 1), then learn the current situation’s model parameters (Lemma 2) and finally predict future transitions (Lemma 3).
@@ IV-A2 Online Situation Modeling:
IV-B Online Situation Identification
To estimate $T_{L}$ (Eq. 9) online we use the Maximum Likelihood Estimate (MLE) of the parameters of $T_{L}$ given the data from the current run $\boldsymbol{x}^{(r)}_{t}$. The estimated parameters are added to a set $\boldsymbol{\Psi}$, where we can keep track of the identified situations. To determine the varying window size of the sequence of past state measurements that make $\boldsymbol{x}^{(r)}_{t}$ we add incoming state estimates to $\boldsymbol{x}^{(r)}_{t}$ until a new changepoint is found. When a new changepoint is found $\boldsymbol{x}^{(r)}_{t}$ is reset to the empty set.
To identify which situation $\psi\in\boldsymbol{\Psi}$ the robot is currently experiencing we employ the average Mahalanobis distance [mahalanobis1930test]:
$M(\psi)\coloneqq\left(\frac{1}{\mid\boldsymbol{x}^{(r)}_{t}\mid}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{x}^{(r)}_{t}}\sqrt{(\mathbf{x}_{i}-\boldsymbol{\mu}_{\psi})^{\top}\Lambda_{\psi}(\mathbf{x}_{i}-\boldsymbol{\mu}_{\psi})}\right),$ (19)
where $\psi=(\boldsymbol{\mu}_{\psi},\Lambda_{\psi})$. In this way we can estimate how far each point in the current run data $\boldsymbol{x}^{(r)}_{t}$ is to the distributions describing the detected local transition models $T_{L}$. Then, using a threshold $\tau$ we determine if the distribution of the closest situation is appropriate to represent the current local dynamics of the robot. If it is not, we create a new situation $\psi=(\boldsymbol{\mu},\Lambda)$ that we add to $\boldsymbol{\Psi}$ during training. At test time, we use the closests experienced situation in $\boldsymbol{\Psi}$ to the one the robot is currently experiencing.
@@ IV-B Online Situation Identification
6:  if $M(\boldsymbol{\mu}_{\psi^{*}},\Lambda_{\psi^{*}})\geq\tau$ and $\mathtt{training}$ then
@@
IV-C Situationally-Aware Dynamics
With the modeled situations representing the local transition dynamics around trajectories in the state space, we proceed to leverage them to obtain the robot’s transition function $T$. As discussed before, we argue that for effective unstructured terrain planning and control we need to account for the hidden state. We do this by representing the hidden latent factors via a parameter $\theta\in\Theta$ that gives $T$ information about the unmodeled world characteristics that also affect the prediction. In such cases, we have a GHP-MDP with transition function:
$T(s,\theta,a,s^{\prime})=\Pr(S_{t+1}=s^{\prime}\mid S_{t}=s,\Theta_{t}=\theta,A_{t}=a),$ (20)
where $S_{t},\Theta_{t},A_{t}\text{ and }S_{t+1}$ are random variables representing the state of the world, the current representation of the latent factors and the actions taken by the agent.
In practical terms, we model the right-hand side of Eq. (20) as an ensemble $\mathcal{M}_{m}:S\times\Theta\times A\rightarrow S$ of $m$ probabilistic (Gaussian) neural networks. The ensemble’s input is augmented with a representation of the current situation (local transition dynamics $T_{L}$) of the robot, in order to inform it about the hidden state of the world. This architecture choice estimates the predictive distribution over the next state and simultaneously considers the epistemic and aleatoric uncertainty [chua2018deep].
IV-C1 Representing the Local Transition Model:
To provide $\mathcal{M}_{m}$ the current situation information we need to map each $\psi\in\boldsymbol{\Psi}$ to a symbol $\theta\in\Theta$. For this, we define the function $\boldsymbol{\Theta}:\boldsymbol{\Psi}\rightarrow\Theta$. In our case, we need $\boldsymbol{\Theta}$ to differentiate between multiple instances of the local distribution $T_{L}$ (Eq. 9), all of which belong to a family of distributions that share a common form but vary in their parameters.
We find that the Moment Generating Function (MGF) provides a useful tool for this purpose, as it uniquely characterizes a distribution [curtiss1942note, Theorem 1]. In general, for a multivariate Gaussian random vector $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu},\Sigma)$, the MGF is defined as
@@ IV-C1 Representing the Local Transition Model:
where we apply the $\log$ function for numerical convenience. Although it could happen that for a given $\boldsymbol{t}$ the representations of two situations would end up being equal, this is highly unlikely and we did not experience any such collisions. Then, whenever $\mathcal{M}_{m}$ is being called, we augment the robot’s state with the current situation’s symbol $\theta$. Note that the current situation is constantly updated as the output from Algorithm 3, which is executed every time a new observation $\boldsymbol{x}_{t}$ is obtained.
IV-C2 Dynamics Learning:
The model $\mathcal{M}_{m}$ is learned by alternating exploration and exploitation, as is common in the MBRL framework. During the first episode, the agent’s actions are considered random as it is relying on a model initialized with random weights that has not yet learned anything about the environment. This exploration phase collects transition information for the replay buffer $\mathcal{R}=\{(s_{i},\theta_{i},a_{i},s_{i+1})\}_{i=1}^{N_{e}}$, where $N_{e}$ is the number of steps in the episode. Note that different from traditional MBRL, we include the representation $\theta_{i}$ of the current situation of the robot in the replay buffer. The collected data is then used to train the dynamics model $\mathcal{M}_{m}$. In subsequent episodes, the agent uses the updated model to make better informed decisions, balancing exploration and exploitation. After each episode, the newly collected data is added to $\mathcal{R}$, and the model is retrained to refine its understanding of the environment’s dynamics. Throughout the training phase, the situation identification process is running in parallel to detect and model the new situations $\psi$, as detailed in Algorithm 3.
Given that the period of time the robot experiences in each situation is different, the replay buffer $\mathcal{R}$ will naturally have an unbalanced sample of situations. Thus, to ensure that the dynamics $\mathcal{M}_{m}$ are learned appropriately for all situations $\psi\in\boldsymbol{\Psi}$, we introduce a loss weighted by the cardinality of each situation. Specifically, we want to minimize the Negative Log-Likelihood (NLL):
IV-C3 Planning and Control:
We use our situationally-aware dynamics model $\mathcal{M}_{m}$ together with the Model Predictive Path Integral (MPPI) method [gandhi2021robust]. The planning process involves generating a set of candidate action sequences over a time horizon $H$ and selecting the sequence that maximizes the expected return. The expected return is computed based on the predicted future states while assuming that the situation symbol $\theta$ remains constant throughout the rollouts. To select an action, MPPI optimizes the function:
$\mathcal{B}(\hat{s}_{t+i})=-\log\|\hat{s}_{t+i}-s_{obs}\|_{2}.$ (26)
Every timestep, the controller applies the first action in $a^{\star}_{t:t+H}$ and then re-evaluates and updates the action sequence, ensuring that the agent adapts to changes in the environment. Figure 2 presents an overview of the system.
V Results

### v3

@@ 4.2 Online and Unsupervised Latent Factor Representation Learning
Our online and unsupervised representation learning method seeks to capture variable and unknown ego- and environmental-factors. This approach models changes in the distribution governing the robot’s transition dynamics as it interacts with the world. These changes encapsulate the influence of latent, unobserved factors that are challenging to predict or exhaustively anticipate. By leveraging this representation of the learned distribution of local dynamics, the robot refines its transition model, enabling more effective planning and control in downstream tasks.
4.2.1 Online Transition Distribution Modeling:
@@
$=\frac{P(S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime})}{P(S_{t}=s,A_{t}=a)}.$ (13)
In particular, $P(s^{\prime}\mid s,a)\propto P(S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime})$. However, as previously noted, in real-world applications, estimates of $P(s^{\prime}\mid s,a)$ usually fail to accurately capture the true dynamics of the world. Generally, this shortfall arises from latent, unobserved factors—such as unmeasured variables or phenomena for which the robot lacks the necessary sensors—that significantly impact state transitions. Thus, our key insight is that by modeling online the joint transition distribution $P(S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime})$ that best explains the UDGP governing the current state transitions, we can account for the latent factors that influence the world dynamics and consequently improve performance on downstream tasks.
To achieve this, we break the problem of online transition distribution modeling into determining local joint transition distributions. By “local”, we mean around the latest set or trajectory of states experienced by the agent. For example, a UGV that is stuck in a slippery area while climbing a mountain will estimate a different local joint transition distribution than the same robot when it is moving over a flat and smooth sidewalk. Figure 1 illustrates this concept, showing how the same action executed from an identical state can lead to vastly different outcomes whenever some important latent factors are not accounted for. Figure 3 demonstrates this effect: it shows marginal transition distributions estimated from real-world UGV data collected while the vehicle executed the same action from comparable initial states. Colored curves correspond to different terrain types, and the dashed black curve denotes the unconditioned (aggregated) estimate, demonstrating how aggregation obscures terrain-specific transition dynamics. [^1] Note that we do not claim that every terrain type requires a separate latent-factor representation. Instead, there is a trade-off between the number of modeled latent factors and their predictive benefit. For example, Figure 21 shows that the distributions for tiles and concrete are very similar, suggesting that a single representation may be sufficient in this case. Crucially, such discrepancies are often better attributed to an incomplete state representation—e.g., the robot lacks sensors, such as a camera, to detect terrain types—rather than inherent aleatoric errors [^2] While our approach does account for aleatoric uncertainty in the transition model, as done in traditional dynamics learning methods, our primary focus is on reducing epistemic uncertainty by identifying and accounting for the latent factors that currently influence the robot’s dynamics..
@@
However, to effectively model the local dynamics using Eq. (14), we need to estimate $\boldsymbol{\eta}=\{\boldsymbol{\mu},\Lambda\}$ online and simultaneously detect when a different $\boldsymbol{\eta}^{\prime}$ describes the current situation of the robot better. For this, we extend BOCD to a multivariate setting and apply it for online and unsupervised discovery of the distinct sets $\boldsymbol{\eta}_{\rho}$. We proceed to introduce our mutivariate extension of BOCD in Section 4.2.2. Then, in Section 4.2.3 we show how we can leverage this results to model online the joint transition distribution of the robot’s dynamics.
4.2.2 Multivariate BOCD:
@@
Since we want to understand when the parameters $\boldsymbol{\eta}$ of the UDGP have changed, we need to define a prior for them. For this, we propose the Normal-Wishart distribution. This conjugate prior is particularly useful because it allows for closed form belief updates for $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and $\Lambda\in\mathbb{R}^{d\times d}$ as more data becomes available. It is defined as follows Murphy (2007):
@@
Essentially, the Normal-Wishart prior combines two distributions: it leverages the Normal distribution for the mean $\boldsymbol{\mu}$, incorporating prior knowledge $\boldsymbol{\mu}_{0}$ about the likely location of $\boldsymbol{\mu}$, and it employs the Wishart distribution to model the precision matrix $\Lambda$, acting as the prior on the inverse covariance. The Wishart distribution is flexible, encoding varying levels of confidence in the scale of the precision matrix through its degrees of freedom parameter and scale matrix, $T$. Parameters $\kappa_{0}$ and $\nu_{0}$ allow us to control the strength of our prior beliefs, with $\kappa_{0}$ scaling the covariance of the mean and $\nu_{0}$ defining the degrees of freedom of the Wishart distribution. This prior choice is satisfactory for our application since it assumes an unknown mean vector $\boldsymbol{\mu}$ and precision matrix $\Lambda$, which enables flexibility by allowing uncertainty in the covariance structure to be encoded in a principled way DeGroot (1970); Murphy (2007). Consequently, it spares us from imposing restrictive assumptions on the covariance matrix, such as isotropy, homogeneity, or mean-dependence, making it highly suitable for modeling multivariate normal distributions with little assumptions about its parameters.
With the sampling model Eq. (15) and prior Eq. (16) determined, we proceed to make the connection to BOCD. Specifically, we will determine the closed form solutions for the growth probability Eq. (5) and changepoint probability Eq. (6), from which we can recover the run length probability Eq. (7). We begin by assuming that the prior on the changepoint probability is $P_{\mathtt{gap}}(g)\sim\mathtt{geometric}(\lambda)$. This makes the process memoryless and the hazard function constant at $H(\tau)=1/\lambda$, where $\lambda$ can be interpreted as the expected time between changepoints Adams and MacKay (2007).
@@
The proof follows from the results of the posterior distribution parameters for the Normal-Wishart prior in (DeGroot, 1970, Sec. 9.10). Specifically, they show that for a set of samples $\{\boldsymbol{x}_{i}:\boldsymbol{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$, the posterior parameters for Eq. (16), given prior parameters $\{\boldsymbol{\mu}_{0},\kappa_{0},T_{0},\nu_{0}\}$, are:
@@
Then, the derivation follows from the fact that $n=1$ for the observation $\boldsymbol{x}_{t}$ at time $t$, and the fact that the parameters $\boldsymbol{\eta}_{t-1}^{(r)}=\{\boldsymbol{\mu}_{t-1}^{(r)},T_{t-1}^{(r)},\nu_{t-1}^{(r)},\kappa_{t-1}^{(r)}\}$ from the previous time step are the current prior. ∎∎
@@
From the BOCD framework, we assumed it is possible to compute the predictive distribution conditional on a given run length Adams and MacKay (2007):
@@
Additionally, $P(\boldsymbol{x}_{t+1}|\boldsymbol{x}_{1:t})$ depends only on the data of the current run $\boldsymbol{x}_{t}^{(r)}$. Thus, $P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{1:t})=P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{t}^{(r)})$. Therefore, using the result from Murphy (2007) that says that the posterior predictive of the Normal-Wishart prior is:
@@
4.2.3 Online Situation Modeling and Identification:
To determine when $T_{L}$ changes and given that we are modeling the system’s situations via distribution parameters $\boldsymbol{\eta}_{0}\coloneqq\{\boldsymbol{\mu}_{0},\kappa_{0},T_{0},\nu_{0}\}$, we can use our previous multivariate extension to BOCD. This process is summarized in Algorithm 2. Concretely, we first determine if we have changed situations (Lemma 1), then learn the current situation’s model parameters (Lemma 2) and finally predict future transitions (Lemma 3).
@@ 4.2.3 Online Situation Modeling and Identification:
To estimate $T_{L}$ (Eq. 14) online we use the Maximum Likelihood Estimate (MLE) of its parameters given the data from the current run $\boldsymbol{x}^{(r)}_{t}$. The estimated parameters are added to our library of situations, $\boldsymbol{\Psi}$. The data for the current run, $\boldsymbol{x}^{(r)}_{t}$, is collected in a variable-sized window; incoming state-action-state tuples are added to $\boldsymbol{x}^{(r)}_{t}$ until a changepoint is detected, at which point $\boldsymbol{x}^{(r)}_{t}$ is reset to the empty set.
To identify which situation $\psi\in\boldsymbol{\Psi}$ the robot is currently experiencing we employ the average Mahalanobis distance Mahalanobis (1930):
$Mah($ $\psi)\coloneqq$ (24)
$\left(\frac{1}{\mid\boldsymbol{x}^{(r)}_{t}\mid}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{x}^{(r)}_{t}}\sqrt{(\mathbf{x}_{i}-\boldsymbol{\mu}_{\psi})^{\top}\Lambda_{\psi}(\mathbf{x}_{i}-\boldsymbol{\mu}_{\psi})}\right),$
where $\psi=(\boldsymbol{\mu}_{\psi},\Lambda_{\psi})$. In this way we can estimate how far each point in the current run data $\boldsymbol{x}^{(r)}_{t}$ is to the distributions describing the detected local transition models. Then, using a threshold $\tau$ we determine if the distribution of the closest situation is appropriate to represent the current local dynamics of the robot. If it is not, we create a new situation $\psi=(\boldsymbol{\mu},\Lambda)$ that we add to $\boldsymbol{\Psi}$ during training. At test time, we use the closest experienced situation in $\boldsymbol{\Psi}$ to the one the robot is currently experiencing. The full identification process is detailed in Algorithm 3.
@@
6:  if $Mah(\psi^{*})\geq\tau$ and $\mathtt{training}$ then
@@
4.2.4 Representing the Local Transition Model:
To provide the dynamics model $\mathcal{M}_{m}$ with the current situation information, we need to map each situation $\psi\in\boldsymbol{\Psi}$ to a symbolic representation $\theta\in\Theta$. For this, we define a function $\boldsymbol{\Theta}:\boldsymbol{\Psi}\rightarrow\Theta$. We need this function to distinguish between multiple instances of the local distribution $T_{L}$ (Eq. 14), all of which belong to a family of distributions that share a common form but vary in their parameters.
We find that the Moment Generating Function (MGF) provides an useful tool for this purpose, as it uniquely characterizes a distribution (Curtiss, 1942, Theorem 1). In general, for a multivariate Gaussian random vector $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu},\Sigma)$, the MGF is defined as
@@ 4.2.4 Representing the Local Transition Model:
where we apply the $\log$ function for numerical convenience. This mapping provides the concrete value for $\theta$ used by the high-level dynamics model in Eq. (8). Although it could happen that for a given $\boldsymbol{t}$ the representations of two situations would end up being equal, this is highly unlikely and we did not experience any such collisions. Then, whenever $\mathcal{M}_{m}$ is being called, we augment the robot’s state with the current situation’s symbol $\theta$. Note that the current situation is constantly updated as the output from Algorithm 3, which is executed every time a new observation $\boldsymbol{x}_{t}$ is obtained.
4.2.5 Additional Details:
To prevent numerical errors while computing the precision matrix during situation creation (Line 7 in Algorithm 3), we also employ an additional hyperparameter describing the minimum support needed. As a rule of thumb, we set it slightly above the dimensionality of the situation space. Additionally, to compute the situation symbol (Eq. 26) we use a constant value of $\mathbf{t}=\mathbf{0.1}$.
5 Results: Unmanned Ground Vehicle